OpenStudy (anonymous):

\[A=\left( \begin{array}{ccc} 2 & -3 & 1 \\ 1 & -2 & 1 \\ 1 & -3 & 2 \end{array} \right)\]

OpenStudy (anonymous):

\[x=\left( \begin{array}{c} 1 \\ 1 \\ 1 \end{array} \right)\]

OpenStudy (anonymous):

\[y=\left( \begin{array}{c} 3 \\ 1 \\ 0 \end{array} \right)\]

OpenStudy (anonymous):

\[z=\left( \begin{array}{c} -1 \\ 0 \\ 1 \end{array} \right)\]

OpenStudy (anonymous):

x , y, z are eigenvector of A

OpenStudy (anonymous):

Find two matrices X and D such that X*D=A*X

OpenStudy (anonymous):

@Zarkon

OpenStudy (zarkon):

this is not unique...choose X to be the zero matrix

OpenStudy (anonymous):

may be they want something other than trivial case?

OpenStudy (amistre64):

D is the diagonal of eugene values; and X is a matrix of columns of (x y z) if those are the eugene vectors

OpenStudy (anonymous):

so how do we find eigenvalue from eigenvector

OpenStudy (amistre64):

\[D=\begin{pmatrix}e_1&0&0\\0&e_2&0\\0&0&e_3 \end{pmatrix}\] \[X=\begin{pmatrix}\vec e_1&\vec e_2&\vec e_3 \end{pmatrix}\]

OpenStudy (amistre64):

the values are unique; just value out A is my first thought

OpenStudy (amistre64):

unique is prolly a bad term

OpenStudy (amistre64):

We could use the definition Ax = Lx , and most times the L is usually a moot point

OpenStudy (amistre64):

i still think Evalueing A is my safest idea

OpenStudy (anonymous):

I tried X D= A X and it doesn't seem to match

OpenStudy (amistre64):

\[det\begin{pmatrix} 2-n & -3 & 1 \\ 1 & -2-n & 1 \\ 1 & -3 & 2-n \end{pmatrix}\] should get us the Evalues

OpenStudy (anonymous):

eigen values :]

OpenStudy (anonymous):

oh,right by finding determinant

OpenStudy (anonymous):

got it , I got \[\lambda1=0,\lambda2=1,\lambda3=1\]

OpenStudy (amistre64):

good, then we rref A by subtracting lambda from the diag to relate it to th vectors

OpenStudy (anonymous):

got it , it matches, thanks so much Amistre

OpenStudy (amistre64):

youre welcome